# How Do LLMs Acquire New Knowledge? A Knowledge Circuits Perspective on Continual Pre-Training

### v1

@@ Abstract
Despite exceptional capabilities in knowledge-intensive tasks, Large Language Models (LLMs) face a critical gap in understanding how they internalize new knowledge, particularly how to structurally embed acquired knowledge in their neural computations. We address this issue through the lens of knowledge circuit evolution, identifying computational subgraphs that facilitate knowledge storage and processing. Our systematic analysis of circuit evolution throughout continual pre-training reveals several key findings: (1) the acquisition of new knowledge is influenced by its relevance to pre-existing knowledge; (2) the evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization; (3) the evolution of knowledge circuits follows a deep-to-shallow pattern. These insights not only advance our theoretical understanding of the mechanisms of new knowledge acquisition in LLMs, but also provide potential implications for improving continual pre-training strategies to enhance model performance [^1] Code and data will be available at https://github.com/zjunlp/DynamicKnowledgeCircuits..
@@ 1 Introduction
Knowledge is a cornerstone of intelligence, shaping how humanity perceives the world, interacts with others, and navigates daily life (Choi, 2022; Chen, 2023). Recent studies (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; DeepSeek-AI et al., 2024; Yang et al., 2024; Zhao et al., 2023; Wu et al., 2024) on Large Language Models (LLMs) have demonstrated their ability to capture factual knowledge from pre-training corpus and encapsulate it as extensive parametric knowledge, empowering their remarkable capabilities in numerous knowledge-intensive tasks (Wang et al., 2024; Cao et al., 2024), as well as in developing higher-order capabilities like reasoning (Qiao et al., 2023; Huang and Chang, 2023). Nevertheless, these powerful models still struggle with knowledge updates, especially with regard to the dynamic nature of world knowledge that evolves after the cut-off date of the pre-training corpus (Zhang et al., 2023; Mousavi et al., 2024). Extensive efforts focus on developing advanced techniques for injecting new knowledge into LLMs (Jang et al., 2022; Jiang et al., 2024; Mecklenburg et al., 2024; Ovadia et al., 2024; Chen et al., 2024a), yet the absence of a well-defined mechanism for new knowledge acquisition in LLMs continues to hinder further progress in this area.
@@
  • 4.2) In the process of knowledge acquisition, the evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization, each marked by unique structural and behaviral characteristics;
@@ 2.1 Circuit Theory
Delving into the Transformer architecture (Vaswani et al., 2017), all computations in a Transformers-based language model as a connected directed acyclic graph, denoted as $\mathcal{G}$. This graph represents the flow of information from the input of the language model to the token unembedding, where activations are projected back to vocabulary space. Various components of a language model, including attention heads and multi-layer perceptrons (MLPs), are defined as the nodes of this graph, denoted as $N$. The edges of this graph, denoted as $E$, are the weighted connections between these components, encompassing residual connections, attention mechanisms, and projections. In the context of Mechanistic Interpretability (MI), which aims to understand the inner workings of adavanced Transformer-based language models (Rai et al., 2024; Ferrando et al., 2024; Bereska and Gavves, 2024; Sharkey et al., 2025), a circuit is conceptualized as a sparse computational subgraph $\mathcal{C}\subset\mathcal{G}$ within a language model whose computations are most relevant to the whole model’s behaviour on the specific task (Olah et al., 2020; Elhage et al., 2021; Wang et al., 2023; Marks et al., 2024). A circuit $\mathcal{C}$ usually contains a selection of nodes $N_{\mathcal{C}}\subset{N}$ and edges $E_{\mathcal{C}}\subset{E}$ necessary for the specific task, expressed as $\mathcal{C}=<N_{\mathcal{C}},E_{\mathcal{C}}>$.
@@ 2.2 Knowledge Circuits
Unlike previous works (Dai et al., 2022; Geva et al., 2021, 2023; Meng et al., 2022) that treat the knowledge blocks as isolated components,  Yao et al. (2024) introduce a novel perspective: knowledge circuits. They hypothesis that the cooperation between multiple components unveils implicit knowledge representation in LLMs. An identified knowledge circuit is considered a computational subgraph that faithfully represents specific knowledge domains within the model’s parametric memory. As such, it should be capable of independently reproducing the behavioral patterns or performance of the entire model with respect to the corresponding tasks. However,  Yao et al. (2024) concentrates exclusively on the knowledge that already stored in the language model, without investigating the process by which LLMs acquire knowledge. In this work, we aim to advance the concept of knowledge circuits by investigating their dynamics throughout continual pre-training.
@@ 4 Analyzing the Evolution of Knowledge Circuits throughout Training
Once we have identified the knowledge circuits, we delve deeper into the changes within the circuits, examining the transitions in the roles and behaviors of nodes and edges. Specifically, we conduct the analysis from three perspectives: performance, topology, and components.
@@ 4.2.2 Topological Centralization
Our results in Figure 3 show a stable downward trend in the knowledge circuit entropy metric for edges in the subgraph across all models, suggesting that the identified knowledge circuits become increasingly centralized, with the importance of critical edges growing as knowledge acquisition progresses. We also observe that the downward trend of the knowledge circuit entropy slows down significantly after a certain turning ponit during the training of all models. For example, turning points are observed in GPT-2 Small, GPT-2 Medium, TinyLlama, and Phi-1.5 at epoch 7, epoch 4, epoch 1, and epoch 1, respectively. We attribute this interesting phenomenon to a phase shift in the evolution of knowledge circuits across continual pre-training. In the initial formation phase of knowledge circuits, less efficient knowledge circuits gradually take shape within the models, resulting in a rapid decrease in circuit entropy. At the phase shift points, the knowledge circuits reach a status of stability where the most critical nodes and edges have been involved. In the subsequent optimization phase, the topology composed critical nodes and edges becomes more stable, while the computations within these components are being optimized to represent and retrieve the knowledge more efficiently, leading to a slowdown in the rate of decrease in circuit entropy.
@@
Takeaway: Biphasic Circuit Evolution The evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization, each marked by unique structural and behaviral characteristics.
This finding suggests that the knowledge circuit state could serve as a valuable tracking status for the continual pre-training process, enabling more informed adjustments to the training method or data in response to different phases. We leave this potential direction for future research.
@@ 4.3.1 Evolutionary Pattern of Components
We first zoom into the specialized nodes within knowledge circuits to investigate the underlying factors driving the evolution of knowledge circuit. Recent studies have identified a set of specialized attention heads (Zheng et al., 2024; Ferrando et al., 2024) that directly contribute to factual recall in Transformer-based LLMs, including the mover head, relation head, and mixture head (Lv et al., 2024; Merullo et al., 2024; Chughtai et al., 2024). More detailed definitions and methodology for identifying these specialized attention heads are provided in Appendix G. We check the emergence and track the proportion of these specialized attention heads in all possible nodes of the knowledge circuits throughout training, and present our results in Figure 5. We observe that during the circuit formation phase, mover heads gradually emerge from nearly zero, while the proportion of relation heads decreases until the phase shift. In the circuit optimization phase, the proportion of all kinds of attention heads stabilizes. The proportion of mixture heads remains stable throughout training. We further examine the layer-wise distribution of mover heads and relation heads within knowledge circuits throuout training. Our results in Figure 6 (and Figure 13 in Appendix) reveal that the increase in mover heads and the decrease in relation heads primarily occur in the mid-to-deeper layers during the circuit formation phase.
@@
Next, we investigate how the nodes within knowledges circuits propagate information to subsequent components through the edges. Specifically, we analyze the variation in edge activation patterns across different layers of the network throughout training. We quantify the edge activation ratio for each layer by calculating the proportion of edges originating from that layer within the knowledge circuit, relative to all possible edges originating from that layer in the whole model [^2] Note that we exclude the activation ratio for the last layer, as the small denominator causes the ratio to be an outlier, potentially blurring the overall trends in the activation patterns observed across layers.. Our results in Figure 7 (and Figure 12 in Appendix) reveal that, during the circuit formation phase, the edges activation ratios in the lower layers gradually decrease, while those in the mid-to-deeper layers exhibit a corresponding increase. However, as training progresses, a transition occurs around the phase shift point, where the edge activation ratios begin to stabilize.
@@
The observed pattern in the evolution of specialized nodes and activated edges within knowledge circuits aligns with the factual recall mechanism in LLMs described by Geva et al. (2023). Specifically, the lower MLP layers specialize in encoding attribute-rich subject representations, while attention heads in the mid-to-deeper layers are responsible for extracting the relevant attributes for a given subject from these lower-level representations. Based on this, we can conclude the evolutionary pattern of knowledge circuits at the component level. During the early training phase of circuit formation, the focus is primarily on developing the extraction function within the nodes of the mid-to-deeper layers of the knowledge circuits. This is reflected in the increased emergence of mover heads and activated edges, along with a decrease in the presence of relation heads in these layers. This process continues until the extraction function is fully established at the phase shift point, as demonstrated by the similar performance advantage of circuits aligned with the post-phase-shift topologies over those aligned with the pre-phase-shift topologies in Figure 4. In the subsequent training phase of circuit optimization, the focus shifts to enriching knowledge representations in the lower layers, evidenced by a stabilized topology and component structure, but with a rapid improvement in the performance of knowledge circuits in Figure 2 and Figure 4.
@@ Appendix A Dataset Construction
Additionally, there are five associated relations—birth date, city, major, university and company—which are randomly sampled from their corresponding pools of possible attributes for each relation. The birthdate relation offers 30 (1 to 30) × 12 (January to December) × 126 (1900 to 2025) possibilities. The corresponding pools of possible attributes for the other four relations are as generated by ChatGPT, as listed in Table 4$\sim$7.
@@ Appendix C Circuit Discovery
Tasks
Unlike previous works that investigate circuits on simple but general tasks such as Indirect Object Identification (IOI) and Greater-Than, our paper focuses on knowledge circuits that are capable of performing the task of factual recall. In a factual recall task, the objective is to predict a target attribute $a$ given a subject-relation pair $(s,r)$. To ensure a sufficiently rich vocabulary space for the first token of the target attribute, we construct the factual recall tasks based on three relations mentioned in §3.1: city, major, and company. The templates for converting a subject-relation pair $(s,r)$ into a query string for each factual recall task are listed in Table 2. A typical circuits task consists of minimal pairs of clean and corrupted inputs. For clean inputs, we randomly sample 300 examples from the training corpus for each knowledge type and frequency as the validation set $D_{\text{val}}$ for circuit discovery. In our experiments, we observe that continually increasing the size of $D_{\text{val}}$ only adds to the runtime for circuit discovery without improving the quality of the discovered circuits. The corresponding corrupted inputs are independently sampled from the training corpus to match the length of the subject tokens in each clean input.
Loss

### v2

@@ Abstract
Despite exceptional capabilities in knowledge-intensive tasks, Large Language Models (LLMs) face a critical gap in understanding how they internalize new knowledge, particularly how to structurally embed acquired knowledge in their neural computations. We address this issue through the lens of knowledge circuit evolution, identifying computational subgraphs that facilitate knowledge storage and processing. Our systematic analysis of circuit evolution throughout continual pre-training reveals several key findings: (1) the acquisition of new knowledge is influenced by its relevance to pre-existing knowledge; (2) the evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization; (3) the evolution of knowledge circuits follows a deep-to-shallow pattern. These insights not only advance our theoretical understanding of the mechanisms of new knowledge acquisition in LLMs, but also provide potential implications for improving continual pre-training strategies to enhance model performance [^1] Code and data are available at https://github.com/zjunlp/DynamicKnowledgeCircuits..
@@ 1 Introduction
Knowledge is a cornerstone of intelligence, shaping how humanity perceives the world, interacts with others, and navigates daily life (Choi, 2022; Chen, 2023). As human society advances, the ways by which knowledge is stored, accessed, and processed have evolved significantly, especailly with the advent of Large Language Models (LLMs). Recent studies (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; DeepSeek-AI et al., 2024; Yang et al., 2024; Zhao et al., 2023; Wu et al., 2024) on LLMs have demonstrated their ability to capture factual knowledge from pre-training corpus and encapsulate it as extensive parametric knowledge, empowering their remarkable capabilities in numerous knowledge-intensive tasks (Wang et al., 2024; Cao et al., 2024), as well as in developing higher-order capabilities like reasoning (Qiao et al., 2023; Huang and Chang, 2023). Nevertheless, these powerful models still struggle with knowledge updates, especially with regard to the dynamic nature of world knowledge that evolves after the cut-off date of the pre-training corpus (Zhang et al., 2023; Mousavi et al., 2024). Extensive efforts focus on developing advanced techniques for injecting new knowledge into LLMs (Jang et al., 2022; Jiang et al., 2024; Mecklenburg et al., 2024; Ovadia et al., 2024; Chen et al., 2024a), yet the absence of a well-defined mechanism for new knowledge acquisition in LLMs continues to hinder further progress in this area.
@@
  • 4.2) In the process of knowledge acquisition, the evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization, each marked by unique structural and behaviral characteristics.
@@ 2.1 Circuit Theory
Delving into the Transformer architecture (Vaswani et al., 2017), all computations in a Transformers-based language model as a connected directed acyclic graph, denoted as $\mathcal{G}$. This graph represents the flow of information from the input of the language model to the token unembedding, where activations are projected back to vocabulary space. Various components of a language model, including attention heads and multi-layer perceptrons (MLPs), are defined as the nodes of this graph, denoted as $N$. The edges of this graph, denoted as $E$, are the weighted connections between these components, encompassing residual connections, attention mechanisms, and projections. In the context of Mechanistic Interpretability (MI), which aims to understand the inner workings of advanced Transformer-based language models (Rai et al., 2024; Ferrando et al., 2024; Bereska and Gavves, 2024; Sharkey et al., 2025), a circuit is conceptualized as a sparse computational subgraph $\mathcal{C}\subset\mathcal{G}$ within a language model whose computations are most relevant to the whole model’s behaviour on the specific task (Olah et al., 2020; Elhage et al., 2021; Wang et al., 2023; Marks et al., 2024). A circuit $\mathcal{C}$ usually contains a selection of nodes $N_{\mathcal{C}}\subset{N}$ and edges $E_{\mathcal{C}}\subset{E}$ necessary for the specific task, expressed as $\mathcal{C}=<N_{\mathcal{C}},E_{\mathcal{C}}>$.
@@ 2.2 Knowledge Circuits
Unlike previous works (Dai et al., 2022; Geva et al., 2021, 2023; Meng et al., 2022) that treat the knowledge blocks as isolated components,  Yao et al. (2024) introduce a novel perspective: knowledge circuits. They hypothesize that the cooperation between multiple components unveils the implicit knowledge representation in LLMs. An identified knowledge circuit is considered a computational subgraph that faithfully represents specific knowledge domains within the model’s parametric memory. As such, it should be capable of independently reproducing the behavioral patterns or performance of the entire model with respect to the corresponding tasks. However,  Yao et al. (2024) concentrates exclusively on the knowledge that already stored in the language model, without investigating the process by which LLMs acquire knowledge. In this work, we aim to advance the concept of knowledge circuits by investigating their dynamics throughout continual pre-training.
@@ 4 Analyzing the Evolution of Knowledge Circuits throughout Training
Once we have identified the knowledge circuits, we delve deeper into the changes within the circuits, examining the transitions in the roles and behaviors of nodes and edges. To improve the clarity and coherence, our analysis follows a three-tiered perspectives, beginning with a surface-level assessment of performance, proceeding to an intermediate exploration of the topology of knowledge circuits, and culminating in a detailed investigation of the underlying components.
@@ 4.2.2 Topological Centralization
Our results in Figure 3 show a stable downward trend in the knowledge circuit entropy metric for edges in the subgraph across all models, suggesting that the identified knowledge circuits become increasingly centralized, with the importance of critical edges growing as knowledge acquisition progresses. We also observe that the downward trend of the knowledge circuit entropy slows down significantly after a certain turning point during the training of all models. For example, turning points are observed in GPT-2 Small, GPT-2 Medium, TinyLlama, and Phi-1.5 at epoch 7, epoch 4, epoch 1, and epoch 1, respectively. We attribute this interesting phenomenon to a phase shift in the evolution of knowledge circuits across continual pre-training. In the initial formation phase of knowledge circuits, less efficient knowledge circuits gradually take shape within the models, resulting in a rapid decrease in circuit entropy. At the phase shift points, the knowledge circuits reach a status of stability where the most critical nodes and edges have been involved. In the subsequent optimization phase, the topology composed critical nodes and edges becomes more stable, while the computations within these components are being optimized to represent and retrieve the knowledge more efficiently, leading to a slowdown in the rate of decrease in circuit entropy.
@@
Moreover, we find that the larger the size of the base pre-trained LLMs, the fewer training steps are required to reach the phase shift point in the knowledge circuits evolution. We suggest that differences in model behavior may stem from the knowledge capacity scaling laws (Allen-Zhu and Li, 2024b), which result from a combination of complex factors such as pre-training data signal-to-noise ratio, pre-training duration and model architectures and warrant further investigation in the future.
Takeaway: Biphasic Circuit Evolution The evolution of knowledge circuits exhibits a distinct phase shift from formation to optimization, each marked by unique structural and behavioral characteristics.
This finding suggests that the state of knowledge circuits could serve as a valuable tracking status for the continual pre-training process, enabling more informed adjustments to the training method or data in response to different phases. We leave this potential direction for future research.
@@ 4.3.1 Evolutionary Pattern of Components
We first zoom into the specialized nodes within knowledge circuits to investigate the underlying factors driving the evolution of knowledge circuit. Recent studies have identified a set of specialized attention heads (Zheng et al., 2024; Ferrando et al., 2024) that directly contribute to factual recall in Transformer-based LLMs, including the mover head, relation head, and mixture head (Lv et al., 2024; Merullo et al., 2024; Chughtai et al., 2024). More detailed definitions and methodology for identifying these specialized attention heads are provided in Appendix G. We check the emergence and track the proportion of these specialized attention heads in all possible nodes of the knowledge circuits throughout training, and present our results in Figure 5. We observe that during the circuit formation phase, mover heads gradually emerge from nearly zero, while the proportion of relation heads decreases until the phase shift. In the circuit optimization phase, the proportion of all kinds of attention heads stabilizes. The proportion of mixture heads remains stable throughout training. We further examine the layer-wise distribution of mover heads and relation heads within knowledge circuits throughout training. Our results in Figure 6 (and Figure 13 in Appendix) reveal that the increase in mover heads and the decrease in relation heads primarily occur in the mid-to-deeper layers during the circuit formation phase.
@@
Next, we investigate how the nodes within knowledge circuits propagate information to subsequent components through the edges. Specifically, we analyze the variation in edge activation patterns across different layers of the network throughout training. We quantify the edge activation ratio for each layer by calculating the proportion of edges originating from that layer within the knowledge circuit, relative to all possible edges originating from that layer in the whole model [^2] Note that we exclude the activation ratio for the last layer, as the small denominator causes the ratio to be an outlier, potentially blurring the overall trends in the activation patterns observed across layers.. Our results in Figure 7 (and Figure 12 in Appendix) reveal that, during the circuit formation phase, the edges activation ratios in the lower layers gradually decrease, while those in the mid-to-deeper layers exhibit a corresponding increase. However, as training progresses, a transition occurs around the phase shift point, where the edge activation ratios begin to stabilize.
@@
The observed pattern in the evolution of specialized nodes and activated edges within knowledge circuits aligns with the factual recall mechanism in LLMs described by Geva et al. (2023). Specifically, the lower MLP layers specialize in encoding attribute-rich subject representations, while attention heads in the mid-to-deeper layers are responsible for extracting the relevant attributes for a given subject from these lower-level representations. Based on this, we can conclude the evolutionary pattern of knowledge circuits at the component level. Since we introduce new knowledge entities via synthetic data that the model did not encounter during pre-training, the extraction function is not yet established for these new knowledge entities at the onset of continual pre-training. Consequently, the model’s attention heads initially concentrate predominantly on the relation tokens already acquired (for example, the city relation learned during pre-training), which manifest as relation heads. During the early training phase of circuit formation, the focus is primarily on developing the extraction function within the nodes of the mid-to-deeper layers of the knowledge circuits. With continual pre-training and the gradual acquisition of new knowledge entities, the attention heads in the model’s mid-to-upper layers increasingly attended to subject tokens, which were thus classified as mover heads. This is reflected in the increased emergence of mover heads and activated edges, along with a decrease in the presence of relation heads in these layers. This process continues until the extraction function is fully established at the phase shift point, as demonstrated by the similar performance advantage of circuits aligned with the post-phase-shift topologies over those aligned with the pre-phase-shift topologies in Figure 4. In the subsequent training phase of circuit optimization, the focus shifts to enriching knowledge representations in the lower layers, evidenced by a stabilized topology and component structure, but with a rapid improvement in the performance of knowledge circuits in Figure 2 and Figure 4.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206246, No. NSFCU23B2055, No. NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023-00138), Yongjiang Talent Introduction Programme (2021A-156-G), CIPSC-SMP-Zhipu Large Model Cross-Disciplinary Fund, Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG, Zhejiang University, the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2023) (Grant A-8001996-00-00). We gratefully acknowledge the support of Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@ Appendix A Dataset Construction
Additionally, there are five associated relations—birth date, city, major, university and company—which are randomly sampled from their corresponding pools of possible attributes for each relation. The birthdate relation offers 30 (1 to 30) × 12 (January to December) × 126 (1900 to 2025) possibilities. The corresponding pools of possible attributes for the other four relations are as generated by ChatGPT, as listed in Table 5$\sim$8.
@@ Appendix C Circuit Discovery
Relation Ratio
birthday 30 / 30351
university 102 / 250
city 151 / 221
major 138 / 188
company 142 / 202
Table 2: Ratio between the unique first tokens and all the possibilities of the attribute.
@@
C.1 Tasks
Unlike previous works that investigate circuits on simple but general tasks such as Indirect Object Identification (IOI) and Greater-Than, our paper focuses on knowledge circuits that are capable of performing the task of factual recall. In a factual recall task, the objective is to predict a target attribute $a$ given a subject-relation pair $(s,r)$. To ensure a sufficiently rich vocabulary space for the first token of the target attribute, we construct the factual recall tasks based on three relations mentioned in §3.1: city, major, and company. We exclude the attributes birthday, whose first token is always an Arabic numeral between 1 and 30, and university, whose first token is typically “University,” from our analysis. We further supplement Table 2 by computing the ratio of unique first tokens to the total number of possible values for each attribute. The findings reveal that the proportion of generic first subtokens is low (approximately 30 %) for the remaining three attributes city, major, and company, thereby mirroring real-world distributions without materially affecting performance evaluation.
The templates for converting a subject-relation pair $(s,r)$ into a query string for each factual recall task are listed in Table 3. A typical circuits task consists of minimal pairs of clean and corrupted inputs. For clean inputs, we randomly sample 300 examples from the training corpus for each knowledge type and frequency as the validation set $D_{\text{val}}$ for circuit discovery. In our experiments, we observe that continually increasing the size of $D_{\text{val}}$ only adds to the runtime for circuit discovery without improving the quality of the discovered circuits. The corresponding corrupted inputs are independently sampled from the training corpus to match the length of the subject tokens in each clean input.
C.2 Loss